During the unidentified process with air, the temperature decreases from 600 0F to 200 0F. Find a. ) the change of internal energy, b.) the change of enthalpy.
1
Expert's answer
2020-05-20T13:50:45-0400

"T_1= 600 ^oF, T_2= 200 ^o F"


Internal Energy = nCvdT= "0.1\times 0.17128 \times (200-600)= -68.512 (BTU)"

for enthalpy of air we can use below formula for finding


"dH=n CpdT=0.1\times 0.24\times (-400)= - 96" BTU